A standardized test's scores are normally distributed with a mean a 500 and a standard deviation of 100. If 1200 students take the test, how many would you expect to score over 650? Round your answer to the nearest whole number.
Answer #1

Let X follows N(u = 500, 0 =100) P(X >650) = P(454, 650 =1-P(z 51.50) Excel function, = 1-0.9332 (= NORMSDIST(-)) = 0.0668 Th